# Empowering recommender systems using automatically generated Knowledge Graphs and Reinforcement Learning

### v1

@@ 4. Experimental Setup
4.1. Automatic KG Generation
To automatically generate KGs from the targeted unstructured data sets, we used two approaches. The first approach makes use of external lexical resources, such as ConceptNet (Speer et al., 2017) to connect terms and enrich the taxonomy. The second approach is different in the way that it neither requires any training nor any external resource, but instead uses the knowledge of the domain available within the input data to extract the knowledge.
4.1.1. ConceptNet-based approach
ConceptNet (Speer et al., 2017) is a knowledge graph that encompasses entities from various domains along with their corresponding relationships. For this study, we specifically focus on three relationship types: IsA, PartOf, and Synonym. The ”IsA” relationship signifies hypernymy relations, while ”PartOf” represents meronymy relations, and ”Synonym” indicates synonymy relations. To generate a dataset for hyponymy relations, we inverted the direction of relations labeled as hypernyms. All other relations in ConceptNet were grouped together as ”other.” The training dataset was created by including all extracted relationships.
The system architecture is based on BERT (Devlin et al., 2019), employing 12 transformer blocks. The embeddings utilized are extracted from the transformer in the 12th layer. Pretrained embeddings from the BERT model ”uncased_L-24_H-1024_A-16” are employed, which are readily available in TensorFlow. We named ”uKG_CN” to the KG that we generated using the ConceptNet-based approach.
4.1.2. Dependency Parsing-based approach
@@
The dependency parsing-based relation extraction approach extracts relations from the text and exports them as triples (left_relation, relation_type, right_relation). It uses dependency parsing (syntactic analysis of the sentences) on the text to find how terms are syntactically (and by extension semantically) connected within sentences. It takes as input the terms extracted by Saffron (Pereira et al., 2019), as well as the dataset originally used to extract the Saffron terms and extract the taxonomy, and returns a list of triples: term1, relation, term2. The whole implementation is done in Python. We named ”uKG_DP” to the KG that we generated using the Dependency Parsing based approach.
We have also created a KG, referred to as ”uKG”, from unstructured data. This KG contains only the article and its relation with the most frequent terms found within the article. To compute the Term Frequency, we utilized TF-IDF.
4.1.3. KG creation using both structured and unstructured data (cKG)
We have already defined the (KGDExR) problem and provided the definition of a KG in section 3. Here, we will illustrate how we constructed KG using both structured and unstructured data (combined data (cKG)). The features of structured data, such as ‘user’, ‘article’, ‘topic’, ‘product’, ‘topic_tag’, ‘product_tag’, ‘response’, etc., serve as the type of nodes or entities in the KG. These entities are connected to other entities through relations such as ‘has_topic’, ‘has_product’, ‘has_topic_tag’, ‘has_product_tag’, and ‘has_response’. Additionally, we utilized the full text of the article, which represents the unstructured data, to create this KG. Therefore, this KG leverages both structured and unstructured data for its creation. The recommendation process begins with a user, traverses through specific entities and their associated relations, and ultimately leads to an item, which in our case is the recommended article for that user. We have named the KG generated using structured and unstructured data that is the combined data as ”cKG”.
4.2. Knowledge Graph Embeddings
In a given KG, each head entity or tail entity can be associated as a point in a continuous vector space. In this work, we use TuckER (Balažević et al., 2019; Arcan et al., 2023) and TransE (Wang et al., 2014) methods to generate KG embeddings. TuckER employs a three-way TuckER tensor decomposition, which computes the tensor T and a sequence of three matrices leveraging the embeddings of entities ($E_{head}$ and $E_{tail}$) and relations ($R$) between them ($G\approx T\otimes E_{head}\otimes R\otimes E_{tail}$).
@@ 4.2. Knowledge Graph Embeddings
4.3. Data sets
The dataset used in this study contains the data of the customers of a large multinational financial services company and the viewpoint articles sent to these customers by the company. The dataset spans from January 30th, 2019 to October 30th, 2019, and contains information of 463 customers who received approximately 80 articles each during this period. The dataset consists of 37,423 rows, detailing individual customer-article interactions. It includes a total of 71 articles, with 66 unique articles, providing details related to the products and services that the financial services company provides. This dataset serves as a valuable resource for researchers and marketers interested in understanding customer behavior and preferences, as well as identifying opportunities for targeted content and marketing strategies. We used this dataset for the evaluation of our KG-driven RL-based approach and KG-driven XGBoost approach for recommending articles to customers. The dataset is divided into training, and test sets with a ratio of 70:30 respectively. We have also made this data set publicly available on a GitHub repository  [^3] https://github.com/GhanshyamVerma/Explainable-Recommender-System..
Table 1. Results of KG-driven XGBoost based Recommender system and KG-driven RL based Recommender system with baseline XGBoost approach.
Model Embedding MAP@K=10 Precision@K=10 Recall@K=10
BPR
(Bayesian personalized ranking)
- 0.11207 0.05672 0.41904
Neighborhood-based
Recommender System
- 0.20418 0.66177 0.27175
NCF
(Neural Collaborative Filtering)
- 0.24104 0.62599 0.30007
XGBoost
Sentence Transformer Embedding
[all-MiniLM-L6-v2]
0.30381 0.65902 0.21568
KG-XGBoost
[ uKG_DP ]
Saffron Dependency Parsing KG
(TuckER KG Embedding)
0.34378 0.71708 0.23965
KG-XGBoost
[ uKG_CN ]
Saffron ConceptNet KG
(TuckER KG Embedding)
0.38985 0.24575 0.74384
KG-XGBoost
[ uKG ]
TransE KG Embedding 0.33774 0.23137 0.70987
KG-XGBoost
[ cKG ]
TransE KG Embedding 0.34468 0.24031 0.73740
KG-RL
[ cKG ]
TransE KG Embedding 0.43761 0.60562 0.24857
5. Results
@@
We constructed two KGs using unstructured data (article text) through Saffron (Pereira et al., 2019) as mentioned in Section 4. These KGs are ”uKG_DP” and ”uKG_CN” where u denotes unstructured data, DP denotes dependency parsing and CN denotes ConceptNet. Additionally, we created a KG referred to as ”cKG” from both structured and unstructured data, as explained in Section 4.
@@
We also compared the performance of our proposed approaches with state-of-the-art existing recommender systems. The existing recommender systems we used are: BPR (Bayesian personalized ranking), Neighborhood-based Recommender System, NCF (Neural Collaborative Filtering), and XGBoost with sentence embedding. We observed that BPR achieved a MAP score of 11.21%, whereas the KG-driven XGBoost approach (cKG) and KG-driven RL-based approach using the same cKG achieved 34.47% and 43.76% MAP scores, respectively. The KG-driven XGBoost approach with KG generated using ConceptNet achieved a MAP score of 38.98% with a recall of 74.38%. The results suggest that if recall is important for any application, then KG-driven XGBoost with uKG_CN can be considered as an option, as it provides the highest recall. Based on the results, it can be observed that the KG-driven RL-based approach outperformed the BPR, Neighborhood-based Recommender System, NCF, and KG-driven XGBoost approaches when considering the MAP score. Additionally, among all the experiments conducted with KG embeddings, the KG embeddings generated from TransE have proven to capture useful information, resulting in better performance compared to TuckER embeddings.
@@
Figure 1. Explaining the recommendations of RL-based approach using the path in the KG that leads to the recommendation.
Explaining the recommendations of RL-based approach using the path in the KG that leads to the recommendation.
Figure 1. Explaining the recommendations of RL-based approach using the path in the KG that leads to the recommendation.
Figure 2. Explaining the recommendations of KG-XGBoost [uKG_CN] model using SHAP.
Figure 3. Explaining the recommendations of KG-XGBoost [uKG_CN] model using ELI5.
@@
Our KG-driven RL-based approach is explainable. To gain a better understanding of our model’s interpretation of the recommendation, we present a case study based on the results obtained from our experiments. In this study, we analyze the path patterns uncovered by our model during the reasoning process, as well as examine different recommendation scenarios. As shown in Figure 1, the article highlighted with a blue dashed boundary is the article recommended by our RL-based model to a user. We can see that the recommended article has some similarities with another article already recommended and clicked by that user, therefore the model thinks that this article should be of relevance for that user as the user was interested in such kind of articles before. Furthermore, our RL-based approach enables us to offer the top 10 articles for each user. Additionally, it can provide all the associated articles in the path that lead to the outcome, along with shared products, topics, and the most frequent common terms found in the text of the articles present in the path. Our RL-based approach can provide such paths for each recommended item to a user which explain the results and play an important role in decision-making.
@@
To generate post-hoc explanation for KG driven XGBoost-based approach, we used SHAP (Lundberg and Lee, 2017) and ELI5444https://github.com/TeamHG-Memex/eli5. SHAP (SHapley Additive exPlanations) is a model-agnsotic method used for explaining the output of machine learning models. It is based on game theoretic concepts and provides an explanation for each feature’s contribution to the model’s prediction. SHAP values quantify the impact of each feature by assigning a value to it, indicating how much it contributes to the prediction compared to the average prediction. SHAP relies on the concept of Shapley values from cooperative game theory and it considers additive feature importance. Figure 2 represents the KG-XGBoost [uKG_CN] model’s features with their average impact on the model output generated by SHAP.
@@
ELI5 (Explain Like I’m 5) is a Python library or framework for explainable machine learning models. ELI5 focuses on understanding the overall behavior and importance of features in making predictions. Eli5 reports feature importance using the ”permutation importance” algorithm. Figure 3 shows the KG-XGBoost [uKG_CN] model’s feature importance by assigning weights to the features based on their impact on the model output generated by ELI5. Both SHAP and ELI5 show that click_frequency, kg_26, article_length, kg_32, Kg_3, and Kg_45 are the most important features that contributed most to the model results.
@@
6. Conclusion
This research paper explores the use of knowledge graphs (KGs) to enhance personalized recommendations in the financial sector. We developed two KG-driven recommender systems for a large multinational financial services company, utilizing reinforcement learning and the XGBoost algorithm, respectively. The first approach employs Reinforcement Learning (RL), while the second utilizes the XGBoost algorithm. The XGBoost-based approach uses KG embeddings generated from both TuckER and TransE, and the RL-based approach uses TransE-generated embeddings. We also performed experiments keeping the KG and the embedding same. The findings suggest that the KG-driven RL-based approach outperforms both the KG-driven XGBoost system and baseline models, delivering more accurate and personalized article recommendations. Additionally, the study emphasizes the importance of reasoning with knowledge for decision-making. Overall, this study highlights the potential of combining advanced machine learning techniques with KG-driven insights to improve customer experience and drive business growth in the investment sector.
Acknowledgements.
This publication has emanated from research supported in part by a grant from Science Foundation Ireland under Grant number SFI/12/RC/2289_P2. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ 4 Experimental Setup
4.1 Automatic KG Generation
To automatically generate KGs from the targeted unstructured data sets, we used two approaches. The first approach makes use of external lexical resources, such as ConceptNet Speer et al. (2017) to connect terms and enrich the taxonomy. The second approach is different in the way that it neither requires any training nor any external resource, but instead uses the knowledge of the domain available within the input data to extract the knowledge.
4.1.1 ConceptNet-based approach
ConceptNet Speer et al. (2017) is a knowledge graph that encompasses entities from various domains along with their corresponding relationships. For this study, we specifically focus on three relationship types: IsA, PartOf, and Synonym. The "IsA" relationship signifies hypernymy relations, while "PartOf" represents meronymy relations, and "Synonym" indicates synonymy relations. To generate a dataset for hyponymy relations, we inverted the direction of relations labeled as hypernyms. All other relations in ConceptNet were grouped together as "other." The training dataset was created by including all extracted relationships.
The system architecture is based on BERT Devlin et al. (2019), employing 12 transformer blocks. The embeddings utilized are extracted from the transformer in the 12th layer. Pretrained embeddings from the BERT model "uncased_L-24_H-1024_A-16" are employed, which are readily available in TensorFlow. We named "uKG_CN" to the KG that we generated using the ConceptNet-based approach.
4.1.2 Dependency Parsing-based approach
@@
The dependency parsing-based relation extraction approach extracts relations from the text and exports them as triples (left_relation, relation_type, right_relation). It uses dependency parsing (syntactic analysis of the sentences) on the text to find how terms are syntactically (and by extension semantically) connected within sentences. It takes as input the terms extracted by Saffron (Pereira et al., 2019), as well as the dataset originally used to extract the Saffron terms and extract the taxonomy, and returns a list of triples: term1, relation, term2. The whole implementation is done in Python. We named "uKG_DP" to the KG that we generated using the Dependency Parsing based approach.
We have also created a KG, referred to as "uKG", from unstructured data. This KG contains only the article and its relation with the most frequent terms found within the article. To compute the Term Frequency, we utilized TF-IDF.
4.1.3 KG creation using both structured and unstructured data (cKG)
We have already defined the (KGDExR) problem and provided the definition of a KG in section 3. Here, we will illustrate how we constructed KG using both structured and unstructured data (combined data (cKG)). The features of structured data, such as ‘subscriber’, ‘educational_article’, ‘topic’, ‘product’, ‘topic_tag’, ‘product_tag’, ‘response’, etc., serve as the type of nodes or entities in the KG. These entities are connected to other entities through relations such as ‘has_topic’, ‘has_product’, ‘has_topic_tag’, ‘has_product_tag’, and ‘has_response’. Additionally, we utilized the full text of the article, which represents the unstructured data, to create this KG. Therefore, this KG leverages both structured and unstructured data for its creation. The recommendation process begins with a subscriber, traverses through specific entities and their associated relations, and ultimately leads to an item, which in our case is the recommended educational article for that subscriber. We have named the KG generated using structured and unstructured data that is the combined data as "cKG".
4.2 Knowledge Graph Embeddings
In a given KG, each head entity or tail entity can be associated as a point in a continuous vector space. In this work, we use TuckER (Balažević et al., 2019; Arcan et al., 2023) and TransE Wang et al. (2014) methods to generate KG embeddings. TuckER employs a three-way TuckER tensor decomposition, which computes the tensor T and a sequence of three matrices leveraging the embeddings of entities ($E_{head}$ and $E_{tail}$) and relations ($R$) between them ($G\approx T\otimes E_{head}\otimes R\otimes E_{tail}$).
@@ 4.2 Knowledge Graph Embeddings
4.3 Data sets
The dataset used in this study contains the data of the subscribers of a large multinational financial services company and the educational articles sent to these subscribers. The dataset spans from January 30th, 2019 to October 30th, 2019, and contains information of 463 subscribers who opted for approximately 80 articles each during this period. The dataset consists of 37,423 rows, detailing individual subscriber-article interactions. It includes a total of 71 educational articles, with 66 unique articles, providing details related to financial products and services. This dataset serves as a valuable resource for researchers and marketers interested in understanding subscriber’s behavior and preferences and choices made by them, as well as identifying opportunities for targeted content and marketing strategies. We used this dataset for the evaluation of our KG-driven RL-based approach and KG-driven XGBoost approach for recommending educational articles to subscribers. The dataset is divided into training, and test sets with a ratio of 70:30 respectively. We have also made this data set publicly available on a GitHub repository  [^3] https://github.com/GhanshyamVerma/Explainable-Recommender-System..
Table 1: Results of KG-driven XGBoost based Recommender system and KG-driven RL based Recommender system with baseline XGBoost approach.
Model Embedding MAP@K=10 Precision@K=10 Recall@K=10 BPR (Bayesian personalized ranking) - 0.11207 0.05672 0.41904 Neighborhood-based Recommender System - 0.20418 0.66177 0.27175 NCF (Neural Collaborative Filtering) - 0.24104 0.62599 0.30007 XGBoost Sentence Transformer Embedding [all-MiniLM-L6-v2] 0.30381 0.65902 0.21568 KG-XGBoost [ uKG_DP ] Saffron Dependency Parsing KG (TuckER KG Embedding) 0.34378 0.71708 0.23965 KG-XGBoost [ uKG_CN ] Saffron ConceptNet KG (TuckER KG Embedding) 0.38985 0.24575 0.74384 KG-XGBoost [ uKG ] TransE KG Embedding 0.33774 0.23137 0.70987 KG-XGBoost [ cKG ] TransE KG Embedding 0.34468 0.24031 0.73740 KG-RL [ cKG ] TransE KG Embedding 0.43761 0.60562 0.24857
5 Results
@@
We constructed two KGs using unstructured data (educational article contents/texts) through Saffron (Pereira et al., 2019) as mentioned in Section 4. These KGs are "uKG_DP" and "uKG_CN" where u denotes unstructured data, DP denotes dependency parsing and CN denotes ConceptNet. Additionally, we created a KG referred to as "cKG" from both structured and unstructured data, as explained in Section 4.
@@
We also compared the performance of our proposed approaches with state-of-the-art (SOTA) existing recommender systems. The existing recommender systems we used are: BPR (Bayesian personalized ranking), Neighborhood-based Recommender System, NCF (Neural Collaborative Filtering), and XGBoost with sentence embedding. We observed that BPR achieved a MAP score of 11.21%, whereas the KG-driven XGBoost approach (cKG) and KG-driven RL-based approach using the same cKG achieved 34.47% and 43.76% MAP scores, respectively. The KG-driven XGBoost approach with KG generated using ConceptNet achieved a MAP score of 38.98% with a recall of 74.38%. The results suggest that if recall is important for any application, then KG-driven XGBoost with uKG_CN can be considered as an option, as it provides the highest recall. Based on the results, it can be observed that the KG-driven RL-based approach outperformed the BPR, Neighborhood-based Recommender System, NCF, and KG-driven XGBoost approaches when considering the MAP score. Additionally, among all the experiments conducted with KG embeddings, the KG embeddings generated from TransE have proven to capture useful information, resulting in better performance compared to TuckER embeddings.
@@
Figure 1: Explaining the recommendations of RL-based approach using the path in the KG that leads to the recommendation.
Figure 2: Explaining the recommendations of KG-XGBoost [uKG_CN] model using SHAP.
@@
Our KG-driven RL-based approach is explainable. To gain a better understanding of our model’s interpretation of the recommendation, we present a case study based on the results obtained from our experiments. In this study, we analyze the path patterns uncovered by our model during the reasoning process, as well as examine different recommendation scenarios. As shown in Figure 1, the educational article highlighted with a blue dashed boundary is the article recommended by our RL-based model to a subscriber. We can see that the recommended article has some similarities with another educational article already recommended and clicked by that subscriber, therefore the model thinks that this article should be of relevance for that subscriber as the subscriber was interested in such kind of articles before. Furthermore, our RL-based approach enables us to offer the top 10 educational articles for each subscriber. Additionally, it can provide all the associated articles in the path that lead to the outcome, along with shared products/services, topics, and the most frequent common terms found in the text of the educational articles present in the path. Our RL-based approach can provide such paths for each recommended item to a user which explain the results and play an important role in decision-making.
Figure 3: Explaining the recommendations of KG-XGBoost [uKG_CN] model using ELI5.
@@
To generate post-hoc explanation for KG driven XGBoost-based approach, we used SHAP Lundberg and Lee (2017) and ELI5444https://github.com/TeamHG-Memex/eli5. SHAP (SHapley Additive exPlanations) is a model-agnostic method used for explaining the output of machine learning models. It is based on game theoretic concepts and provides an explanation for each feature’s contribution to the model’s prediction. SHAP values quantify the impact of each feature by assigning a value to it, indicating how much it contributes to the prediction compared to the average prediction. SHAP relies on the concept of Shapley values from cooperative game theory and it considers additive feature importance. Figure 2 represents the KG-XGBoost [uKG_CN] model’s features with their average impact on the model output generated by SHAP.
@@
ELI5 (Explain Like I’m 5) is a Python library or framework for explainable machine learning models. ELI5 focuses on understanding the overall behavior and importance of features in making predictions. Eli5 reports feature importance using the "permutation importance" algorithm. Figure 3 shows the KG-XGBoost [uKG_CN] model’s feature importance by assigning weights to the features based on their impact on the model output generated by ELI5. Both SHAP and ELI5 show that click_frequency, kg_26, article_length, kg_32, Kg_3, and Kg_45 are the most important features that contributed most to the model results.
@@
6 Conclusion
This research paper explores and demonstrates the use of knowledge graphs (KGs) to enhance personalized recommendations in the financial sector. We proposed two KG-driven recommender systems for a large multinational financial services company, utilizing reinforcement learning and the XGBoost algorithm, respectively. The first approach employs Reinforcement Learning (RL), while the second utilizes the XGBoost algorithm. The XGBoost-based approach uses KG embeddings generated from both TuckER and TransE, and the RL-based approach uses TransE-generated embeddings. We also performed experiments keeping the KG and the embedding same. The findings suggest that the KG-driven RL-based approach outperforms both the KG-driven XGBoost system and baseline models, delivering more accurate and personalized educational article recommendations. Additionally, the study emphasizes the importance of reasoning with knowledge for decision-making. Overall, this study highlights the potential of combining advanced machine learning techniques with KG-driven insights to improve customer experience and drive business growth in the investment sector.
Acknowledgments
This publication has emanated from research supported in part by a grant from Science Foundation Ireland under Grant number SFI/12/RC/2289_P2 {Insight} and a grant from Fidelity Investments. For the purpose of Open Access, the authors have applied a CC BY public copyright license to any Author Accepted Manuscript version arising from this submission.